OpenStudy (anonymous):

x+6y=2 (1) x=5-6y (2) what is the ordered pair for this system of equations

OpenStudy (anonymous):

x+6y=2 . x=5-6y rewrite x+6y=5

OpenStudy (anonymous):

time (-1): x+6y=2 u get -x-6y=-2 -x-6y=-2 x+6y=5 Add : ----------- 0= 3 no solution

OpenStudy (anonymous):

click good answer if I deserve a medal